OpenStudy (anonymous):

A small dog is trained to jump straight up a distance of 1.2 m. How much kinetic energy does the 7.2-kg dog need to jump this high? (The acceleration due to gravity is 9.8 m/s2.) Show your work.

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

I am sorry....I am not the best at physics

OpenStudy (rajat97):

okay this is a good question here, you need to use the kinematical equation and the kinetic energy both collectively interesting.... so first of all let us analyze the things that are given to us so first of all the vertical displacement s(displacement)=1.2m mass of the dog=7.2kg=m acceleration = -9.8m/s^2(-ve sign as the acceleration is in the downward direction) and the final velocity=v=0(at maximum height, the direction of the motion of the dog will be reversed so the final velocity at the topmost point will be zero) so first of all , we will apply the kin. equation v^2 - u^2=2as where v=final velocity and u=initial velocity and a=acceleration and s is the displacement so by putting the above values into the equation, we get 0-u^2=2 x 9.8 x 1.2 so u^2=23.52 now , we can find the kinetic energ by the formula K.E.=(mv^2)/2 so by putting the values in this equation, we get K.E.=84.627J well, this all was just useless(if you know the work energy theorem):) you can do it by the wok energy theorem as follows the work energy theorem says that, total work done=change kinetic energy of the body we know that final kinetic energy = 0(initial vel.=0) and only one force i.e. gravity is acting on the body so the work done by gravity on the body=change in its kinetic energy -mgh=(mv^2)/2-(mu^2)/2 (-mgh as force is actinginthe opposite direction of the displacement) here,v=0 so -mgh=-(mu^2)/2 so, mgh=(mu^2) so by putting the values of m,g and h, you'll get the kinetic energy required and it will be the same as before hope this helps you:)